OpenStudy (selfconfidence777):

-x-y=6 a. -1,-4 b. -2,-3 c. -4,-2 d. -3,-2

OpenStudy (igreen):

What are you trying to find?

OpenStudy (selfconfidence777):

which order pair is the solution of this equation?

OpenStudy (selfconfidence777):

@iGreen

OpenStudy (igreen):

Oh, just plug in the points and see if they satisfy the equation. (-1, -4) -x - y = 6 1 + 4 = 6

OpenStudy (igreen):

1 + 4 = ?

OpenStudy (selfconfidence777):

5

OpenStudy (selfconfidence777):

doesnt it have to equal to 6

OpenStudy (igreen):

Yes, so we have: 5 = 6 This is incorrect..so (-1, -4) isn't a solution to the equation.

OpenStudy (selfconfidence777):

so which one is?

OpenStudy (igreen):

Plug in (-4, -2) -x - y = 6 4 + 2 = 6 What's 4 + 2?

OpenStudy (selfconfidence777):

6

OpenStudy (selfconfidence777):

lol thanks